# Impurity immersed in a two-component few-fermion mixture in a one-dimensional harmonic trap

### v2

@@ I Introduction
In this work, we analyze a mesoscopic fermionic mixture consisting of three distinguishable components. Our aim is to test whether any collective effects may appear in the case of a larger number of components, which are not present in binary mixtures. In order to make these studies well-defined, reasonably comprehensive, and also going beyond extensively studied balanced mixtures, we consider the smallest possible extension of a two-component fermionic mixture by introducing additional interaction with a single third-component particle. We aim to investigate if and how interactions with this impurity change the internal properties of the mixture. Moreover, we inspect non-classical correlations this may induce. In this way, we also link our work to the well-studied problem of a single impurity interacting with several spin-polarized fermions Wenz et al. (2013); Włodzyński (2022).
Our work is organized as follows. In Sec. II we introduce the theoretical model of the system studied and we describe its Hamiltonian in the convenient formalism. Then in Sec. III we briefly explain the exact diagonalization method in the energy cut-off framework. In Sec. IV we analyze single- and two-particle density distributions and we identify a specific transition in the many-body ground state of the system induced by interactions with impurity. In Sec. V we study this transition in detail and we show that it has intriguing universal properties that do not depend essentially on either the number of particles in the system or the internal interactions in the mixture. Finally, we conclude in Sec. VI.
@@ II The system
where the dimensionless parameter $g$ is the effective interaction strength. Note that we assumed here that interactions with the third-component particle are the same for both components. Generalization to non-symmetric interactions is straightforward but beyond the scope of this work. We treat both interaction strengths $g_{0}$ and $g$ as independent parameters. It is justified from an experimental point since in principle they depend differently on the external magnetic field Chin et al. (2010), the shape of the trapping potential in the transverse direction Olshanii (1998), and the channel of interactions determined by the total spin of colliding particles Xu et al. (2017).
@@
Let us also underline that in our approach we intentionally use different formalisms for different components, i.e., the two-component mixture is described in the second quantization which automatically takes into account the fermionic statistics, while the third-component particle is described within the first quantization emphasizing that there is only one particle there. In this notation, the total Hamiltonian $\hat{\cal H}$ acts in the Hilbert space being a tensor product of a single-particle Hilbert space of the third particle and a Fock space in the occupation number representation of the two-component mixture. This formalism has previously been used successfully to describe spinless fermions interacting with single impurity Włodzyński (2022).
@@
In the Hamiltonian (1) we already assumed many limiting simplifications: one-dimensionality, an equal mass of all particles, the same parabolic trap, zero-range $s$-wave interactions, or symmetric interaction with the third component. Still, however, there is a huge flexibility related to the number of particles in components $A$ and $B$. Fortunately, the Hamiltonian (1) commutes independently with operators of the total number of particles $\hat{N}_{\sigma}=\int\!\mathrm{d}x\,\hat{\Psi}^{\dagger}_{\sigma}(x)\hat{\Psi}_{\sigma}(x)$. It means that its properties can be studied independently in the subspaces of a well-defined number of particles. It is known that in the case of two-component mixtures (in our case equivalent to setting $g=0$) measurable properties of the system strongly depend on the particle-number imbalance. This is related to a specific competition between the external potential constantly compressing the cloud of all particles, the quantum statistics acting inside fermionic components acting as quantum repealing pressure, and the inter-component interactions tending to spatially separate different components. To make our analysis as clear as possible, in this work, we focus only on symmetric mixtures having the same number of particles in both components, $N=N_{A}=N_{B}$. This assumption, together with the assumption of symmetric interactions with the third-component particle, leads directly to a full symmetry in the mixture sector. This symmetry is manifested formally by the fact that any eigenstate of the many-body Hamiltonian (1) is invariant under the interchange of A and B components.
@@ III The Method
To find the many-body ground state of the Hamiltonian (1) we adopt the numerically exact diagonalization technique. To make it feasible and accurate we need to build an appropriately tuned many-body basis in which the Hamiltonian is represented and diagonalized. Since all particles are confined in the same external parabolic potential, it is convenient to choose the eigenstates of the harmonic oscillator as a single-particle basis for all components. They have a known textbook form
@@
Of course, they fulfill natural anticommutation fermionic relations. This gives us a natural path to introduce basis in the many-body Hilbert space. As mentioned already, in our notation, the basis vectors have a form of tensor product
@@
where $|\phi_{n}\rangle$ is a state of the third-component particle occupying $n$-th single-particle orbital, while two integer vectors $\vec{\alpha}=(\alpha_{1},\ldots,\alpha_{N})$ and $\vec{\beta}=(\beta_{1},\ldots,\beta_{N})$ encode single-particle orbitals occupied by individual fermions from components $A$ and $B$, respectively. Indistinguishability and fermionic nature of particles is guaranteed by imposing additional constrains of the form $\alpha_{1}>\alpha_{2}>\ldots>\alpha_{N}$ and $\beta_{1}>\beta_{2}>\ldots>\beta_{N}$. States (5) are obviously eigenstates of the non-interacting Hamiltonian ($g_{0}=g=0$) and have energies $E_{\vec{\alpha},\vec{\beta},n}=\hbar\Omega\left[(2N+1)/2+n+\sum_{i}(\alpha_{i}+\beta_{i})\right]$. Non-interacting ground state has energy $E_{0}=\hbar\Omega(N^{2}+1/2)$.
@@
To diagonalize numerically the Hamiltonian one needs to cut the basis. In our approach, we do not perform simple cuts on single-particle excitations but rather we limit the many-body basis to states having total (non-interacting) energy $E_{\vec{\alpha}\vec{\beta}n}$ not larger than some fixed energy limit $E_{C}$. In this way, we do not exclude from the description low-energy states having only a few highly excited particles, while we neglect high-energy states with many low-excited particles. It was shown in previous studies that such a method gives a much better approximation of interacting many-body states with the same usage of numerical resources Haugset and Haugerud (1998); Płodzień et al. (2018); Chrostowski and Sowiński (2019); Rojo-Francàs et al. (2022). In our case, determination of the appropriate value of cut-off energies $E_{C}$ for specific systems considered is based on a convergence test of the many-body ground state. A detailed description of this procedure and the cut-off energy values used is provided in the Appendix A. Following this procedure, for a given number of particles $N$ and interaction strengths $g_{0}$ and $g$, we represent the Hamiltonian (1) in the truncated many-body basis and diagonalize with the implicitly-restarted Arnoldi-Lanczos algorithm Lehoucq et al. (1998). As a result, we obtain the lowest many-body eigenstates represented as a superposition of basis states $|{\vec{\alpha},\vec{\beta},n}\rangle$ and their energies. In this way, we have particular access to a whole structure of the many-body ground state $|\mathtt{G}\rangle$. We present exemplary spectra for the system of five particles ($N=2$) as a function of the interaction strengths in Fig. 1. Let us mention here that in the case of harmonic confinement, further optimization techniques of the many-body basis are possible Volosniev, A.G. et al. (2015); Bellotti, Filipe F. et al. (2017); Kościk (2018, 2020). They are vital when large interaction strengths are considered. In the cases of relatively weak interactions studied in our work, this kind of optimization has limited importance. Therefore, we do not introduce them to our scheme.
@@ IV Impact of the impurity
Figure 2: Impact of interactions with the third-component particle on the two-component mixture containing four particles ($N=2$). Different columns correspond to different values of interaction strength with the impurity $g$. In all plots, inter-component interaction strength is fixed, $g_{0}=1$. (a) Single-particle density profiles of the mixture’s components $n_{\sigma}(x)$ (thick blue) and the third-component particle $m(x)$ (thin red). For a better comparison, the non-interacting distribution of the mixture’s components is additionally displayed with a light blue shadow. (b) Single-particle reduced density matrix of the impurity $\rho(x,x^{\prime})$. (c) Two-particle density profile $\mu_{\sigma}(x,z)$. (d) Distribution of correlations ${\cal G}_{\sigma}(x,z)$ between impurity and particle from the component $\sigma$. Note, that in the left plot, we consider the interaction strength $g$ slightly positive ($g=0.01$) since for vanishing interactions the distribution vanishes. In this case, the distribution intensity is also scaled by a factor $10^{2}$ to make it visible. The upper and lower values on the color bar refer to plots (c) and (d), respectively. In all plots, positions are given in units of $a_{0}$. In panels (a) and (b), densities are expressed in units of $a_{0}^{-1}$, while in panels (c)-(d), in units of $a_{0}^{-2}$.
Figure 3: Distribution of internal two-particle correlations ${\cal K}(x,y)$ in the system containing impurity interacting with the two-component mixture containing four particles ($N=2$) for different internal interaction strengths $g_{0}$ and interactions with the impurity $g$. For any fixed interaction $g_{0}$, along with increasing repulsion $g$, the transition between two different regimes is visible. In all plots, positions are expressed in units of $a_{0}$ while densities in units of $a_{0}^{-2}$.
In principle, the impact of interactions with an additional impurity on the system can be analyzed from two complementary perspectives, depending if one is interested in the external or internal behavior of the mixture. First, let us discuss the external features when interactions with the impurity are present. The simplest quantities that feel an influence of additional interaction are single-particle density distributions $n_{\sigma}(x)$ of mixture components and $m(x)$ of the impurity. They are straightforwardly defined as
@@
where $\hat{\mathbb{P}}_{x}=|x\rangle\langle x|$ projects to an eigenstate $|x\rangle$ of position operator. Of course, due to the exchange symmetry $A\leftrightarrow B$ explained before, both density distributions $n_{A}(x)$ and $n_{B}(x)$ are exactly the same. In Fig. 2(a), we present these distributions (thick blue and thin red lines) for the system of five particles ($N=2$) for different interaction strengths $g$, keeping fixed internal interaction $g_{0}=1$. For better comparison, we additionally display with the shadowed area the non-interacting ($g_{0}=g=0$) distribution $n_{\sigma}(x)$. It is clear that along with increasing $g$, the distribution of the impurity $m(x)$ significantly changes its character and for sufficiently large repulsions it is split and pushed out from the center where the mixture is concentrated. In contrast, distributions of mixture components $n_{\sigma}(x)$ are only slightly affected. Since the impact of interactions on the impurity is significant, it is also worth analyzing the off-diagonal order of the separated particle. This is encoded directly in the single-particle reduced density matrix which in our notation can be written as
@@
Obviously, a diagonal part of this quantity is equal to a single-particle density, $m(x)=\rho(x,x)$. As clearly seen in Fig. 2(b), the reduced density matrix undergoes a substantial change when the system is driven between the two regimes.
@@
Up to now, we presented illustrative results for a system containing five particles (two particles in each mixture’s component and the impurity). For completeness, we performed analogous calculations for systems with a larger number of particles in the mixture (up to $N=5$). They are exposed in the Supplemental Material Teske and Sowiński (2025). On a quantitative level, the results obtained are very similar in the sense that the observed structural change in the many-body ground state is universal. Namely, for a given number of particles and fixed internal interaction $g_{0}$, one finds such interaction strength $g$ at which distributions change rapidly. Particularly, the single-particle density $m(x)$ is split and pushed to the edges while the distribution of internal correlations ${\cal K}(x,y)$ changes its nature.
@@ V The structural transition
As explained already, the system studied experiences some substantial transition in its many-body ground state, and a precise quantitative study of its nature is needed. To get a better understanding of this transition we utilize additional quantitative tools. Since the method of exact diagonalization gives direct access to the structure of the many-body ground state one can calculate the fidelity between ground states obtained for different interaction strengths. When internal interaction $g_{0}$ is fixed, the fidelity is defined as
@@
In this expansion $\chi(g_{0};g)=\partial^{2}_{g^{\prime}}{\cal F}(g_{0};g,g^{\prime})|_{g^{\prime}=g}$ is the fidelity susceptibility which quantifies the resistivity of the ground state on infinitesimal change of interaction strength You et al. (2007); Wang et al. (2015). Its large amplitude for a given value of $g$ signals that at this point the ground state of the system rapidly changes with $g$, i.e., the state of the system quickly moves toward different regions of the Hilbert space. In our numerical approach, we find susceptibility $\chi(g_{0};g)$ straightforwardly by using a three-point approximation of a second derivative with a small step $\delta g=g^{\prime}-g=10^{-4}$.
@@
In principle, the $N$-dependent universal curve $\phi_{N}(\xi)$ is different for different numbers of particles in the mixture. For example, it is visible that for a larger number of particles, it becomes significantly narrower, indicating that the transition is more rapid. We found, however additional simple scaling between systems of different numbers of particles which makes all the curves reasonably overlapping. Namely, all $N$-dependent functions $\phi_{N}(\xi)$ can be fused to a single universal curve $\Phi(\xi)$ by applying the scaling $\phi_{N}(\xi)=\Phi(N^{2}\xi)$. In Fig. 5(d), we display all the fidelity susceptibilities from panel (a) after the aforementioned scaling and shifting. We see that all the transition plots follow almost the same curve, and the agreement is better for a larger number of particles. Consequently, we argue that any fidelity susceptibility curve can be derived straightforwardly from the universal function $\Phi(\xi)$ via the relation

### v3

@@ I Introduction
In this work, we analyze a mesoscopic fermionic mixture consisting of three distinguishable components. Our aim is to test whether any collective effects may appear in the case of a larger number of components, which are not present in binary mixtures. In order to make these studies well-defined, reasonably comprehensive, and also going beyond extensively studied balanced mixtures, we consider the smallest possible extension of a two-component fermionic mixture by introducing additional interaction with a single third-component particle. We aim to investigate if and how interactions with this impurity change the internal properties of the mixture. Moreover, we inspect non-classical correlations that this may induce. In this way, we also link our work to the well-studied problem of a single impurity interacting with several spin-polarized fermions Wenz et al. (2013); Włodzyński (2022).
Our work is organized as follows. In Sec. II we introduce the theoretical model of the system studied and we describe its Hamiltonian in the convenient formalism. Then, in Sec. III we briefly explain the exact diagonalization method in the energy cut-off framework. In Sec. IV we analyze single- and two-particle density distributions and we identify a specific transition in the many-body ground state of the system induced by interactions with the impurity. In Sec. V we study this transition in detail and we show that it has intriguing universal properties that do not depend essentially on either the number of particles in the system or the internal interactions in the mixture. Finally, we conclude in Sec. VI.
@@ II The system
where the dimensionless parameter $g$ is the effective interaction strength. Note that we assumed here that interactions with the third-component particle are the same for both components. Generalization to non-symmetric interactions is straightforward but beyond the scope of this work. We treat both interaction strengths $g_{0}$ and $g$ as independent parameters. It is justified from an experimental point since, in principle, they depend differently on the external magnetic field Chin et al. (2010), the shape of the trapping potential in the transverse direction Olshanii (1998), and the channel of interactions determined by the total spin of colliding particles Xu et al. (2017).
@@
Let us also underline that in our approach we intentionally use different formalisms for different components, i.e., the two-component mixture is described in the second quantization, which automatically takes into account the fermionic statistics, while the third-component particle is described within the first quantization emphasizing that there is only one particle there. In this notation, the total Hamiltonian $\hat{\cal H}$ acts in the Hilbert space being a tensor product of a single-particle Hilbert space of the third particle and a Fock space in the occupation number representation of the two-component mixture. This formalism has previously been used successfully to describe spinless fermions interacting with single impurity Włodzyński (2022).
@@
In the Hamiltonian (1) we already assumed many limiting simplifications: one-dimensionality, an equal mass of all particles, the same parabolic trap, zero-range $s$-wave interactions, or symmetric interaction with the third component. Still, however, there is a huge flexibility related to the number of particles in components $A$ and $B$. Fortunately, the Hamiltonian (1) commutes independently with operators of the total number of particles $\hat{N}_{\sigma}=\int\!\mathrm{d}x\,\hat{\Psi}^{\dagger}_{\sigma}(x)\hat{\Psi}_{\sigma}(x)$. It means that its properties can be studied independently in the subspaces of a well-defined number of particles. It is known that in the case of two-component mixtures (in our case, equivalent to setting $g=0$), measurable properties of the system strongly depend on the particle-number imbalance. This is related to a specific competition between the external potential constantly compressing the cloud of all particles, the quantum statistics acting inside fermionic components acting as quantum repelling pressure, and the inter-component interactions tending to spatially separate different components. To make our analysis as clear as possible, in this work, we focus only on symmetric mixtures having the same number of particles in both components, $N=N_{A}=N_{B}$. This assumption, together with the assumption of symmetric interactions with the third-component particle, leads directly to a full symmetry in the mixture sector. This symmetry is manifested formally by the fact that any eigenstate of the many-body Hamiltonian (1) is invariant under the interchange of A and B components.
@@ III The Method
To find the many-body ground state of the Hamiltonian (1) we adopt the numerically exact diagonalization technique. To make it feasible and accurate, we need to build an appropriately tuned many-body basis in which the Hamiltonian is represented and diagonalized. Since all particles are confined in the same external parabolic potential, it is convenient to choose the eigenstates of the harmonic oscillator as a single-particle basis for all components. They have a known textbook form
@@
Of course, they fulfill natural anticommutation fermionic relations. This gives us a natural path to introduce a basis in the many-body Hilbert space. As mentioned already, in our notation, the basis vectors have the form of a tensor product
@@
where $|\phi_{n}\rangle$ is a state of the third-component particle occupying $n$-th single-particle orbital, while two integer vectors $\vec{\alpha}=(\alpha_{1},\ldots,\alpha_{N})$ and $\vec{\beta}=(\beta_{1},\ldots,\beta_{N})$ encode single-particle orbitals occupied by individual fermions from components $A$ and $B$, respectively. Indistinguishability and fermionic nature of particles are guaranteed by imposing additional constraints of the form $\alpha_{1}>\alpha_{2}>\ldots>\alpha_{N}$ and $\beta_{1}>\beta_{2}>\ldots>\beta_{N}$. States (5) are obviously eigenstates of the non-interacting Hamiltonian ($g_{0}=g=0$) and have energies $E_{\vec{\alpha},\vec{\beta},n}=\hbar\Omega\left[(2N+1)/2+n+\sum_{i}(\alpha_{i}+\beta_{i})\right]$. The non-interacting ground state has energy $E_{0}=\hbar\Omega(N^{2}+1/2)$.
@@
To diagonalize the Hamiltonian numerically, one needs to cut the basis. In our approach, we do not perform simple cuts on single-particle excitations, but rather we limit the many-body basis to states having total (non-interacting) energy $E_{\vec{\alpha}\vec{\beta}n}$ not larger than some fixed energy limit $E_{C}$. In this way, we do not exclude from the description low-energy states having only a few highly excited particles, while we neglect high-energy states with many low-excited particles. It was shown in previous studies that such a method gives a much better approximation of interacting many-body states with the same usage of numerical resources Haugset and Haugerud (1998); Płodzień et al. (2018); Chrostowski and Sowiński (2019); Rojo-Francàs et al. (2022). In our case, the determination of the appropriate value of cut-off energies $E_{C}$ for specific systems considered is based on a convergence test of the many-body ground state. A detailed description of this procedure and the cut-off energy values used is provided in the Appendix A. Following this procedure, for a given number of particles $N$ and interaction strengths $g_{0}$ and $g$, we represent the Hamiltonian (1) in the truncated many-body basis and diagonalize with the implicitly-restarted Arnoldi-Lanczos algorithm Lehoucq et al. (1998). As a result, we obtain the lowest many-body eigenstates represented as a superposition of basis states $|{\vec{\alpha},\vec{\beta},n}\rangle$ and their energies. In this way, we have particular access to a whole structure of the many-body ground state $|\mathtt{G}\rangle$. We present exemplary spectra for the system of five particles ($N=2$) as a function of the interaction strengths in Fig. 1. Firstly, one notes that the spectrum appears to display features that could be indicative of quantum chaos, although confirming this would require more detailed analysis. This type of analysis goes significantly beyond the scope of this work, and we leave it for future studies. Importantly for our studies, we note that the ground state is isolated for any interaction parameters $g_{0}$ and $g$, and degeneracy may occur only in the limit of infinite repulsions. This is a general observation valid for any number of particles we consider, and it means that any observed structural changes in the ground state do not occur due to the intersection of states in the spectrum and are an inherent property of the ground state itself.
Let us mention here that in the case of harmonic confinement, further optimization techniques of the many-body basis are possible Volosniev, A.G. et al. (2015); Bellotti, Filipe F. et al. (2017); Kościk (2018, 2020). They are vital when large interaction strengths are considered. In the cases of relatively weak interactions studied in our work, this kind of optimization has limited importance. Therefore, we do not introduce them to our scheme.
@@ IV Impact of the impurity
Figure 2: Impact of interactions with the third-component particle on the two-component mixture containing four particles ($N=2$). Different columns correspond to different values of interaction strength with the impurity $g$. In all plots, inter-component interaction strength is fixed, $g_{0}=1$. (a) Single-particle density profiles of the mixture’s components $n_{\sigma}(x)$ (thick blue) and the third-component particle $m(x)$ (thin red). For a better comparison, the non-interacting distribution of the mixture’s components is additionally displayed with a light blue shadow. (b) Single-particle reduced density matrix of the impurity $\rho(x,x^{\prime})$. (c) Two-particle density profile $\mu_{\sigma}(x,z)$. (d) Distribution of correlations ${\cal G}_{\sigma}(x,z)$ between impurity and particle from the component $\sigma$. Note that in the left plot, we consider the interaction strength $g$ slightly positive ($g=0.01$) since for vanishing interactions the distribution vanishes. In this case, the distribution intensity is also scaled by a factor $10^{2}$ to make it visible. The upper and lower values on the color bar refer to plots (c) and (d), respectively. In all plots, positions are given in units of $a_{0}$. In panels (a) and (b), densities are expressed in units of $a_{0}^{-1}$, while in panels (c)-(d), in units of $a_{0}^{-2}$.
Figure 3: Distribution of internal two-particle correlations ${\cal K}(x,y)$ in the system containing impurity interacting with the two-component mixture containing four particles ($N=2$) for different internal interaction strengths $g_{0}$ and interactions with the impurity $g$. For any fixed interaction $g_{0}$, along with increasing repulsion $g$, the transition between two different regimes is visible. In all plots, positions are expressed in units of $a_{0}$ while densities are in units of $a_{0}^{-2}$.
In principle, the impact of interactions with an additional impurity on the system can be analyzed from two complementary perspectives, depending on whether one is interested in the external or internal behavior of the mixture. First, let us discuss the external features when interactions with the impurity are present. The simplest quantities that feel an influence of additional interaction are single-particle density distributions $n_{\sigma}(x)$ of mixture components and $m(x)$ of the impurity. They are straightforwardly defined as
@@
where $\hat{\mathbb{P}}_{x}=|x\rangle\langle x|$ projects to an eigenstate $|x\rangle$ of position operator. Of course, due to the exchange symmetry $A\leftrightarrow B$ explained before, both density distributions $n_{A}(x)$ and $n_{B}(x)$ are exactly the same. In Fig. 2(a), we present these distributions (thick blue and thin red lines) for the system of five particles ($N=2$) for different interaction strengths $g$, keeping fixed internal interaction $g_{0}=1$. For better comparison, we additionally display with the shadowed area the non-interacting ($g_{0}=g=0$) distribution $n_{\sigma}(x)$. It is clear that along with increasing $g$, the distribution of the impurity $m(x)$ significantly changes its character and for sufficiently large repulsions it is split and pushed out from the center where the mixture is concentrated. In contrast, distributions of mixture components $n_{\sigma}(x)$ are only slightly affected. Since the impact of interactions on the impurity is significant, it is also worth analyzing the off-diagonal order of the separated particle. This is encoded directly in the single-particle reduced density matrix, which in our notation can be written as
@@
Obviously, a diagonal part of this quantity is equal to a single-particle density, $m(x)=\rho(x,x)$. As clearly seen in Fig. 2(b), the reduced density matrix undergoes a substantial change when the system is driven between the two regimes. In particular, distinct isolated enhancements appear at the end of the diagonal and anti-diagonal. They are a direct consequence of splitting and pulling apart the probability distribution of a single particle.
@@
Up to now, we have presented illustrative results for a system containing five particles (two particles in each mixture’s component and the impurity). For completeness, we performed analogous calculations for systems with a larger number of particles in the mixture (up to $N=5$). They are exposed in the Supplemental Material Teske and Sowiński (2025). On a quantitative level, the results obtained are very similar in the sense that the observed structural change in the many-body ground state is universal. Namely, for a given number of particles and fixed internal interaction $g_{0}$, one finds such an interaction strength $g$ at which distributions change rapidly. Particularly, the single-particle density $m(x)$ is split and pushed to the edges while the distribution of internal correlations ${\cal K}(x,y)$ changes its nature.
@@ V The structural transition
As explained already, the system studied experiences some substantial transition in its many-body ground state, and a precise quantitative study of its nature is needed. To get a better understanding of this transition, we utilize additional quantitative tools. Since the method of exact diagonalization gives direct access to the structure of the many-body ground state, one can calculate the fidelity between ground states obtained for different interaction strengths. When internal interaction $g_{0}$ is fixed, the fidelity is defined as
@@
In this expansion, $\chi(g_{0};g)=\partial^{2}_{g^{\prime}}{\cal F}(g_{0};g,g^{\prime})|_{g^{\prime}=g}$ is the fidelity susceptibility which quantifies the resistivity of the ground state on infinitesimal change of interaction strength You et al. (2007); Wang et al. (2015). Its large amplitude for a given value of $g$ signals that at this point the ground state of the system rapidly changes with $g$, i.e., the state of the system quickly moves toward different regions of the Hilbert space. In our numerical approach, we find susceptibility $\chi(g_{0};g)$ straightforwardly by using a three-point approximation of a second derivative with a small step $\delta g=g^{\prime}-g=10^{-4}$.
@@
In principle, the $N$-dependent universal curve $\phi_{N}(\xi)$ is different for different numbers of particles in the mixture. For example, it is visible that for a larger number of particles, it becomes significantly narrower, indicating that the transition is more rapid. We found, however, additional simple scaling between systems of different numbers of particles, which makes all the curves reasonably overlapping. Namely, all $N$-dependent functions $\phi_{N}(\xi)$ can be fused to a single universal curve $\Phi(\xi)$ by applying the scaling $\phi_{N}(\xi)=\Phi(N^{2}\xi)$. In Fig. 5(d), we display all the fidelity susceptibilities from panel (a) after the aforementioned scaling and shifting. We see that all the transition plots follow almost the same curve, and the agreement is better for a larger number of particles. Consequently, we argue that any fidelity susceptibility curve can be derived straightforwardly from the universal function $\Phi(\xi)$ via the relation